Fermi National Laboratory
 
Operation Summary for February 17, 2012 - February 20, 2012
 

NuMI reported receiving 3.67E18 protons on target for the period from 2/13/12 to 2/20/12.

Friday February 17

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI and MiniBooNE experiments taking beam.

During the day shift, Operators prepared enclosure C for a 20-minute access at 8:37 AM. Operators searched and secured the Meson MTest area for the EDIT School and began sending them beam at 11:24 AM. Booster and MI personnel began a phase-lock study at 12:38 PM. A Booster person began a study at 3:37 PM. The Booster MI phase-lock study ended at 3:43 PM.

During the evening shift, a Linac RF station (LRF5) tripped off; it reset. The MTest EDIT School stopped using beam for the day at 5:34 PM. Operations established beam to the Meson FTBF experiment T-979 at 6 PM.

Saturday February 18

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, Operations disabled beam to T-979 at 12:17 AM. A Linac Klystron RF station tripped off at 3:51 AM; it reset. Operations reported at 5:14 AM that the MiniBooNE beam permit had tripped off seven times during the shift due to bad beam positions; Operators made adjustments.

During the day shift, at 9:27 AM, Operations reported that three ion pumps in the MI RF sector tripped off; they reset. Operations established beam to T-979 at 10:12 AM. At 12:21 PM, a Booster RF station (BRF1)1 tripped off; it reset. Water Group personnel reported at 12:23 PM that the CUB column #1 had completed its regeneration; they reported that it would be ready for operation in about 24 hours. Operators completed a MI tune change at 12:31 PM. Operators searched and secured the Neutrino Area N01 enclosure and turned on power supplies at 3:54 PM.

During the evening shift, Operations disabled beam to T-979 at 4:51 PM; this ended their current run, but they might return for more beam within a week. At 6:49 PM, Operators completed a 750 KeV line and 400 MeV area tune. A Booster vacuum pump (BIPRF22) tripped off and brought down BRF11 and 12; everything recovered and reset. At 9:58 PM, MIRF16 tripped off; it reset. At 11:05 PM, the I- Source sparked and its output dropped; Operators adjusted it back to normal.

Sunday February 19

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, at 12:40 AM, Operators reported there had been 13 I- Source sparks in the last hour with Operators adjusting it back to its normal output. At 1:38 AM, MI lost its ramp due to a MI-30 quadrupole magnet ground fault indication; everything reset. Operators reported at 5:13 AM that for the last 40 minutes the I- Source sparks had become more frequent; Operations called in Linac personnel. At 7:05 AM, the NuMI beam permit tripped off on an incomplete MI RF gap indication; it reset. At 7:59 AM, the MI South Anode Power Supply tripped off on an interlock fault; it reset.

During the day shift, a Linac person reported at 8:05 AM that they had to switch to the H- Source; Operations disabled the Linac beam. The Linac beam returned at 8:26 AM. Operators completed a tune of the 750 KeV line at 12:32 PM. Operators completed a Booster injection tune at 2:16 PM. At 3:48 PM, LRF2 tripped off; it reset.

During the evening shift, at 4:02 PM, BRF3 tripped off; it reset. At 5:38 PM, MIRF10 tripped off; it reset. At 6:22 PM, BRF7 and KRF3 tripped off; they both reset.

Monday February 20

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, at 3:53 AM, KRF1 tripped off; it reset. At 3:57AM, KRF3 tripped off; it reset. Operators searched and secured the Neutrino Area NM3 and NM4 enclosures; there were some problems with NM3 that Operators will report to Interlock personnel when the show up in the morning.

Operations Coordinator (OpsCo)

The OpsCo wants the complex to operate with as few beam disruptions as possible for the EDIT School and the other experiments.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/20/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory